# SpikeCLIP: A Contrastive Language-Image Pretrained Spiking Neural Network

### v3

@@ 4.2 Zero-shot Results
Table 4: Comparison of energy consumption on six image classification benchmarks. The application of SpikeCLIP results in an average energy reduction of approximately $78\%$.
@@
4.3 Ablation Study
In order to assess the impact of knowledge distillation during pre-training and the introduction of the KL-divergence term during the fine-tuning phase on SpikeCLIP’s performance, we conducted a series of ablation studies across six distinct datasets. These studies involved eliminating the pre-training stage and removing the KL-divergence term from the loss function. As indicated by the results presented in Table 3, the full-fledged SpikeCLIP consistently achieved the highest accuracy across all evaluated datasets. Moreover, the incorporation of KD-based pre-training and the KL-divergence term enhanced the average performance by $22.45\%$ and $4.92\%$, respectively. These findings indicate the critical role of pre-training via knowledge distillation in enabling SpikeCLIP to deliver performance on par with its ANN counterparts, and the use of KL-divergence significantly boosts SpikeCLIP’s performance in a variety of image classification tasks.
(a) Accuracy on CIFAR10 dataset.
(b) Accuracy on ImageNet-1k dataset.
We would like to understand how the size and data distributions of the pre-training datasets affect SpikeCLIP’s performance. In addition to investigating the impact of diverse dataset sizes, we explored the effect of the degree of overlap between pre-training data and downstream task data on performance (as shown in Figure 4). We evaluated three different data distributions during the pre-training phase: one included one-third of the downstream task data (indicated by “More similar”), another excluded downstream task data completely (indicated by “Less similar”), and the third represented a scenario falling between these two extremes (indicated by “No similarity”). For further details on the creation of different pre-training datasets, please refer to Appendix E. As hypothesized, a direct correlation is observed between the growth in the size of pre-training data and the usage of more downstream task data during the pre-training phase. These experimental findings suggest the potential for further enhancing SpikeCLIP’s performance by expanding the size and coverage of the pre-training dataset.
4.4 Comparison of Energy Consumption
@@
5 Conclusion
We found it hard to train SNNs for multimodal tasks directly due to the challenge of integrating linguistic and visual features into a unified representation through spike trains. To circumvent this obstacle, we suggested a two-step training recipe: an initial phase of pre-training for cross-modal alignment via knowledge distillation, followed by dual-loss fine-tuning using surrogate gradients. A readout mechanism was proposed to interpret the states of SNNs for enabling knowledge distillation from ANNs, and a regularization term was introduced to preserve the generalization capacity attained during the pre-training phase. Through extensive experimentation on $6$ image classification datasets, we demonstrated that the SNNs trained with the proposed method can match the performance of their ANN counterparts in multimodal classification tasks and exhibit zero-shot learning capabilities.
Limitations
In this study, we have embarked on one of the initial endeavors to employ spiking neural networks (SNNs) in multimodal tasks, with a particular emphasis on classification tasks. It would indeed be intriguing to broaden the scope of this work to encompass generative tasks, such as image captioning and visual question answering. We relied on an existing CLIP to pre-train SpikeCLIP by distilling knowledge from the conventional CLIP. In the future, we plan to collect a larger collection of image-text pairs, and it would be interesting to explore whether we can pre-train SNNs in an unsupervised manner from scratch, using the collected image-text pairs. This could potentially enable us to narrow the performance disparity between SNNs and cutting-edge ANN models.
@@ 5 Conclusion
The authors have made great efforts to ensure the reproducibility of the empirical results reported in this paper. To begin with, the experiment settings, evaluation metrics, and datasets were described in detail in Subsections 4.1, 4.2, and Appendix B, E, G. Furthermore, the implementation details were clearly presented in Subsections 3.4, 3.5 and Appendix A, C. In our effort to ensure reproducibility, we have submitted the source code of the proposed training algorithm with our paper, and plan to release the source code on GitHub upon acceptance.
Appendix
@@ Appendix A Implementation Details of Training Method
For the pre-trained CLIP model, we use openai/clip-vit-base-patch16 with a dimension of 512 in this study. A Spikingformer-4-384 ([Zhou et al., 2023b]) with 4 layers and a dimension of 384 is used as the base model for comparison. The image-side component architecture of SpikeCLIP is built upon this base model with a time-step weight (TSW) layer followed by a dimensionality-mapping layer, aligning the output to a 512-dimensional space compatible with pre-trained CLIP models.
For comparing purposes with SpikeCLIP, we constructed ScratchCLIP as an ANN counterpart. The image encoder of ScratchCLIP has a 4-layer transformer and uses a patch-splitting layer with the same number of parameters as SpikeCLIP. ScratchCLIP’s text encoder uses an MLP architecture, as well as a word embedding layer of conventional CLIP.
@@
    • ImageNet-1k: The ImageNet-1k serves as a foundational benchmark in computer vision research, comprising approximately 1.2 million high-resolution color images across 1,000 distinct categories. The dataset is commonly partitioned into training, validation, and testing subsets to enable rigorous evaluation of machine learning models. Due to its scale and diversity, ImageNet-1k has become instrumental in the development and assessment of state-of-the-art algorithms. In addition, this dataset is one of the largest image classification datasets available[Russakovsky et al., 2015].
@@
    • CIFAR10: The CIFAR10 serves as a well-established benchmark within the domains of machine learning and computer vision. Comprising 60,000 color images with a resolution of 32x32 pixels, the dataset is organized into 10 unique classes. With each class containing 6,000 images, the dataset ensures a balanced class distribution. Conventionally, CIFAR10 is partitioned into 50,000 images for training and 10,000 images for testing, thereby providing a consistent framework for evaluating the performance of classification models[Krizhevsky, 2009].
@@
    • CIFAR100: An extension of the CIFAR10 dataset, CIFAR100 is also a prominent benchmark in the fields of machine learning and computer vision. While maintaining the same overall count of 60,000 color images at a 32x32 pixel resolution, CIFAR100 expands the class diversity to 100 distinct categories, each represented by 600 images. For evaluative purposes, the dataset is typically segmented into 50,000 training images and 10,000 testing images. This augmented class variety enhances CIFAR100’s utility for conducting more nuanced assessments of classification models[Krizhevsky, 2009].
@@
    • Flower102: The Flower102 dataset is a notable asset within the computer vision landscape, explicitly designed to cater to fine-grained image recognition endeavors. The dataset comprises a diverse set of images, capturing 102 different floral species. Each category is scrupulously curated to maintain a balanced representation, thereby enabling more sophisticated model evaluations. Due to its focus on capturing subtle variances between closely aligned classes, the Flower102 dataset plays a pivotal role in both refining and benchmarking specialized image classification algorithms[Nilsback and Zisserman, 2008].
@@
    • Caltech101: As an esteemed benchmark in computer vision research, the Caltech101 dataset encompasses an assemblage of approximately 9,000 color images, categorized into 101 distinct object classes. These classes span a diverse array of subjects, including animals, vehicles, and inanimate objects, with a fluctuating number of images allocated to each category. Widely employed for a variety of computational tasks, such as object recognition and classification, Caltech101 offers a multifaceted visual dataset for the rigorous evaluation of machine learning model performance[Fei-Fei et al., 2004].
@@
    • OxfordIIIPet: The OxfordIIIPet dataset holds a significant position in the realm of computer vision, particularly in the context of fine-grained classification assignments. The dataset comprises visual representations of 37 distinct breeds of cats and dogs, furnishing a nuanced foundation for algorithms engineered to discern subtle visual cues. Each breed category is populated with a balanced assortment of images, thereby facilitating the compilation of representative training and testing subsets. Owing to its targeted emphasis on the classification of pet breeds, the OxfordIIIPet dataset proves invaluable for fine-tuning models aimed at specialized image recognition tasks[Parkhi et al., 2012].
@@
    Appendix C Impact of Learnable Time-dependent Weights on Spiking Integration
    In previous SNNs, tensor values were averaged across different time steps ($T$) before being classified. However, this approach assigns the same weight to each step ($1/T$), ignoring their interdependence. In particular, if the previous time step has already produced a spike, it may be more difficult for the current time step to produce a new spike again, so the signal from the new spike generated by the current time step may be stronger. This idea is not considered in cases where different time steps are given the same weight, which can lead to reduced performance.To address this issue, we employ learnable parameters to replace the fixed averaging weights, which are incorporated into SpikeCLIP.
      Dataset Baseline AD AR TDW
    CIFAR10 $94.39$ $94.39$ $94.45$ $\bm{94.48}$
    CIFAR100 $77.51$ $77.58$ $77.56$ $\bm{77.69}$
    Appendix D Impact of Text Encoder Architectures on SpikeCLIP’s Performance
    An improvement was noted upon switching to a Multi-Layer Perceptron (MLP-based) architecture for the text encoder, by following the work Bal and Sengupta [2023]. Our observation suggested that within the two-step training scheme (Pre-training + Fine-tuning), the text encoder is prone to overfitting if the architecture is overly complex and the architecture of MLP is proven to be proficient in this study. Comprehensive experimental results are presented in Table.6.
    Transformer-based $86.37$ $48.03$ $75.78$ $27.09$ $33.93$ $94.76$ $60.99$
    MLP-based $90.63$ $64.69$ $79.88$ $62.86$ $81.79$ $97.58$ $79.57$
    The LayerNorm layer is used in the text information processing of this study since it is indispensable in the text data processing. However, the hardware chip design should be based on the design of relevant algorithms. The wide application of the BatchNorm layer in SNN architecture is inseparable from the algorithm research of SNN in image processing.
    Appendix E Pre-training Dataset Sizes and Distributions
@@
    Appendix F Designing More Challenging Multimodal Image Classification Tasks
@@
    The selection process was facilitated through a specific prompt, termed Prompt1, with the assistance of ChatGPT [OpenAI, 2022]. Additionally, we conducted four sub-experiments involving random label replacement at different scales: 20%, 40%, 80%, and 100%. For the initial three scenarios, predefined random seeds were used, and each was executed in triplicate to record both the mean and variance of the results.
@@
    Appendix G Comparison of Energy Consumption
    According to Yao et al. [2022], the theoretical energy consumption of layer $l$ in a SNN can be calculated as:
@@ Appendix G Comparison of Energy Consumption
    where FLOPs is the floating point operations of $b$, which is the number of multiply-and-accumulate (MAC) operations. Assuming that the MAC and AC operations are implemented on the 45nm hardware [Horowitz, 2014], where $E_{MAC}=4.6pJ$ and $E_{AC}=0.9pJ$ ($1$J $=10^{3}$ mJ $=10^{12}$ PJ).
@@
    Therefore, we estimate the theoretical energy consumption of SpikeCLIP as follows:
@@
    where SNN FC and SNN Conv are the fully connected linear layer and the convolutional layer with neurons in SpikeCLIP respectively. As shown in Equation.10, the SOPs of $m$ SNN Fully Connected Layer (FC), $n$ SNN Convolutional layers are added together and multiplied by $E_{AC}$.
    We refer to Horowitz [2014], assuming that MAC and AC operations are implemented on 45nm hardware (the calculation of power consumption in this hardware only involves MAC and AC operations) since SpikeCLIP and ScratchCLIP have the same architecture except for pulsar neurons, We can calculate the energy consumption reduction (ECR) by equations 7, 8, 9 and 10 as the following expression equation:

### v4

@@ 4.3 Ablation Study
As indicated by the results presented in Table 3, the full-fledged SpikeCLIP consistently achieved the highest accuracy across all evaluated datasets. Moreover, the incorporation of KD-based pre-training and the KL-divergence term enhanced the average performance by $22.45\%$ and $4.92\%$, respectively. These findings indicate the critical role of pre-training via knowledge distillation in enabling SpikeCLIP to deliver performance on par with its ANN counterparts, and the use of KL-divergence significantly boosts SpikeCLIP’s performance in a variety of image classification tasks.
Table 4: Performance on CIFAR10 and Flowers102 datasets for different $\lambda_{1}$ values. Setting 1 refers to the scenario where, after fine-tuning in the second stage, the model is evaluated on the same dataset used for fine-tuning. Setting 2 refers to the case where, after fine-tuning on one dataset, the model is evaluated on the other dataset.
$\lambda_{1}$ Setting 1 Setting 2
CIFAR10 Flowers102 CIFAR10 Flowers102
$0.5$ $94.48$ $86.88$ $91.25$ $74.52$
$1.0$ $94.48$ $86.17$ $90.85$ $74.23$
$2.0$ $94.45$ $86.25$ $91.05$ $74.15$
$10.0$ $94.27$ $86.08$ $90.44$ $73.52$
Secondly, we also explore the impact of the KL regularization term’s coefficient $\lambda_{1}$ in the second stage. We conducted the hyper-parameter sensitive analysis of $\lambda_{1}$, as shown in Table 4. We perform experiments in two settings: Setting 1 refers to the scenario where the second-stage fine-tuning is performed on a specific dataset, and the evaluation is carried out on the same dataset. On the other hand, Setting 2 refers to the scenario where the second-stage fine-tuning is performed on one dataset, but evaluation is done on a different dataset. We found that the value of $\lambda_{1}$ has minimal impact on the performance of SpikeCLIP during the second stage.
Finally, we aim to investigate how the size and data distributions of the pre-training datasets influence SpikeCLIP’s performance. In addition to investigating the impact of diverse dataset sizes, we also explored how the degree of overlap between pre-training data and downstream task data influences performance (Figure 6). We evaluated three different data distributions during the pre-training phase: one included one-third of the downstream task data (indicated by “More similar”), another excluded downstream task data completely (indicated by “Less similar”), and the third represented a scenario falling between these two extremes (indicated by “No similarity”). Further details on the creation of these datasets are provided in C. As hypothesized, a direct correlation is observed between the growth in the size of pre-training data and the usage of more downstream task data during the pre-training phase. These experimental findings suggest the potential for further enhancing SpikeCLIP’s performance by expanding the size and coverage of the pre-training dataset.
(a) Accuracy on CIFAR10 dataset.
(b) Accuracy on ImageNet-1k dataset.
4.4 Impact of Text Encoder Architectures
Transformer-based $86.37$ $48.03$ $75.78$ $27.09$ $33.93$ $94.76$ $60.99$
MLP-based $90.63$ $64.69$ $79.88$ $62.86$ $81.79$ $97.58$ $79.57$
An improvement was noted upon switching to a Multi-Layer Perceptron (MLP-based) architecture for the text encoder, by following the work [41]. Our observation suggested that within the two-step training scheme (Pre-training + Fine-tuning), the text encoder is prone to overfitting if the architecture is overly complex and the architecture of MLP is proven to be proficient in this study. Comprehensive experimental results are presented in Table 5.
4.5 Impact of Learnable Time-dependent Weights on Spiking Integration
In previous SNNs, tensor values were averaged across different time steps ($T$) before being classified. However, this approach assigns the same weight to each step ($1/T$), ignoring their interdependence. In particular, if the previous time step has already produced a spike, it may be more difficult for the current time step to produce a new spike again, so the signal from the new spike generated by the current time step may be stronger.
This idea is not considered in cases where different time steps are given the same weight, which can lead to reduced performance. To address this issue, we employ learnable parameters to replace the fixed averaging weights, which are incorporated into SpikeCLIP.
  Dataset Baseline AD AR TDW
CIFAR10 $94.39$ $94.39$ $94.45$ $\bm{94.48}$
CIFAR100 $77.51$ $77.58$ $77.56$ $\bm{77.69}$
4.6 Comparison of Computing Energy Consumption
Table 7: Estimation of computing energy consumption on six image classification benchmarks. The application of SpikeCLIP results in an average energy reduction of approximately $78\%$.
@@ 4.6 Comparison of Computing Energy Consumption
Most importantly, we would like to clarify that our energy consumption estimation focuses solely on computing energy, excluding factors such as memory access and data movement. A detailed discussion of this limitation can be found in G.
5 Conclusion and Future Work
Conclusion  We found it hard to train SNNs for multimodal tasks directly due to the challenge of integrating linguistic and visual features into a unified representation through spike trains. To circumvent this obstacle, we suggested a two-step training recipe: an initial phase of pre-training for cross-modal alignment via knowledge distillation, followed by dual-loss fine-tuning using surrogate gradients. A readout mechanism was proposed to interpret the states of SNNs to enable knowledge distillation from ANNs, and a regularization term was introduced to preserve the generalization capacity attained during the pre-training phase. Through extensive experimentation on $6$ image classification datasets, we demonstrated that the SNNs trained with the proposed method can match the performance of their ANN counterparts in multimodal classification tasks and exhibit zero-shot learning capabilities.
Future Work  The following are our plans for scaling up to larger datasets in future work to improve generalization capabilities, particularly for zero-shot tasks: First, it is indeed challenging to obtain a dataset of the same scale as CLIP’s full training dataset, as you said. However, with the rapid development of large multimodal models, we anticipate that access to larger synthetic multimodal datasets, such as LAION-5B [66], will become feasible in the future. Then, we will follow the experimental settings of CLIP [19] and conduct zero-shot experiments after the pre-training on the large datasets. Secondly, the primary objective of this work was to explore the feasibility of modality fusion between text and images using a spiking neural network (SNN) architecture. Our experiments demonstrate that this approach is viable and promising, providing a potential pathway to reduce the energy consumption of future multimodal large models. Finally, we note that in biological systems, multimodal signals (such as sound, images, and speech) are processed using spike signals. Our work validates the biological plausibility of integrating multimodal information using spiking neural networks. We believe that these insights can contribute to the future development of more efficient multimodal models. Limitations are discussed in G.
@@ 5 Conclusion and Future Work
The authors have made great efforts to ensure the reproducibility of the empirical results reported in this paper. To begin with, the experiment settings, evaluation metrics, and datasets were described in detail in Section 4.1, Section 4.2, and B, C, F. Furthermore, the implementation details were clearly presented in Section 3.5, Section 3.6 and A, Section 4.5. In our effort to ensure reproducibility, we have submitted the source code of the proposed training algorithm with our paper, and plan to release the source code on GitHub upon acceptance.
Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable comments. This work was supported by National Natural Science Foundation of China (No. 62076068).
@@ Appendix A Implementation Details of Training Method
For the pre-trained CLIP model, we use openai/clip-vit-base-patch16 with a dimension of 512 in this study. A Spikingformer-4-384 ([15]) with 4 layers and a dimension of 384 is used as the base model for comparison. The image-side component architecture of SpikeCLIP is built upon this base model with a time-step weight (TSW) layer followed by a dimensionality-mapping layer, aligning the output to a 512-dimensional space compatible with pre-trained CLIP models.
For comparing purposes with SpikeCLIP, we constructed ScratchCLIP as an ANN counterpart. The image encoder of ScratchCLIP has a 4-layer Transformer and uses a patch-splitting layer with the same number of parameters as SpikeCLIP.
ScratchCLIP’s text encoder uses an MLP architecture, as well as a word embedding layer of conventional CLIP.
@@
      ImageNet-1k: The ImageNet-1k serves as a foundational benchmark in computer vision research, comprising approximately 1.2 million high-resolution color images across 1,000 distinct categories. The dataset is commonly partitioned into training, validation, and testing subsets to enable rigorous evaluation of machine learning models. Due to its scale and diversity, ImageNet-1k has become instrumental in the development and assessment of state-of-the-art algorithms. In addition, this dataset is one of the largest image classification datasets available[51].
    • 2.
@@
      CIFAR10: The CIFAR10 serves as a well-established benchmark within the domains of machine learning and computer vision. Comprising 60,000 color images with a resolution of 32x32 pixels, the dataset is organized into 10 unique classes. With each class containing 6,000 images, the dataset ensures a balanced class distribution. Conventionally, CIFAR10 is partitioned into 50,000 images for training and 10,000 images for testing, thereby providing a consistent framework for evaluating the performance of classification models[58].
    • 3.
@@
      CIFAR100: An extension of the CIFAR10 dataset, CIFAR100 is also a prominent benchmark in the fields of machine learning and computer vision. While maintaining the same overall count of 60,000 color images at a 32x32 pixel resolution, CIFAR100 expands the class diversity to 100 distinct categories, each represented by 600 images. For evaluative purposes, the dataset is typically segmented into 50,000 training images and 10,000 testing images. This augmented class variety enhances CIFAR100’s utility for conducting more nuanced assessments of classification models[58].
    • 4.
@@
      Flower102: The Flower102 dataset is a notable asset within the computer vision landscape, explicitly designed to cater to fine-grained image recognition endeavors. The dataset comprises a diverse set of images, capturing 102 different floral species. Each category is scrupulously curated to maintain a balanced representation, thereby enabling more sophisticated model evaluations. Due to its focus on capturing subtle variances between closely aligned classes, the Flower102 dataset plays a pivotal role in both refining and benchmarking specialized image classification algorithms[64].
    • 5.
@@
      Caltech101: As an esteemed benchmark in computer vision research, the Caltech101 dataset encompasses an assemblage of approximately 9,000 color images, categorized into 101 distinct object classes. These classes span a diverse array of subjects, including animals, vehicles, and inanimate objects, with a fluctuating number of images allocated to each category. Widely employed for a variety of computational tasks, such as object recognition and classification, Caltech101 offers a multifaceted visual dataset for the rigorous evaluation of machine learning model performance[61].
    • 6.
@@
      OxfordIIIPet: The OxfordIIIPet dataset holds a significant position in the realm of computer vision, particularly in the context of fine-grained classification assignments. The dataset comprises visual representations of 37 distinct breeds of cats and dogs, furnishing a nuanced foundation for algorithms engineered to discern subtle visual cues. Each breed category is populated with a balanced assortment of images, thereby facilitating the compilation of representative training and testing subsets. Owing to its targeted emphasis on the classification of pet breeds, the OxfordIIIPet dataset proves invaluable for fine-tuning models aimed at specialized image recognition tasks[62].
    • 7.
@@
    Appendix C Pre-training Dataset Sizes and Distributions
@@
    Appendix D Designing More Challenging Multimodal Image Classification Tasks
@@
    The selection process was facilitated through a specific prompt, termed Prompt1, with the assistance of ChatGPT [67]. Additionally, we conducted four sub-experiments involving random label replacement at different scales: 20%, 40%, 80%, and 100%. For the initial three scenarios, predefined random seeds were used, and each was executed in triplicate to record both the mean and variance of the results.
@@
    Appendix E Results of Zero-shot Experiments of SpikeCLIP on 26 Datasets
    we follow the setup of CLIP [19] to conduct zero-shot experiments using $26$ datasets. We exclude ImageNet since it has been used for pre-training. The results are presented in the Table 8.
    Table 8: We demonstrated the zero-shot generalization ability of ScratchCLIP and SpikeCLP on 26 datasets.
    Model
    FER2013
    STL10
    EuroSAT
    RESISC45
    GTSRB
    KITTI
    Country211
    PCAM
    UCF101
    Kinetics700
    CLEVR
    HatefulMemes
    SST
    StratchCLIP $47.20$ $75.69$ $68.42$ $55.40$ $52.95$ $31.80$ $2.70$ $21.58$ $9.80$ $7.50$ $2.18$ $24.74$ $36.40$
    SpikeCLIP $45.72$ $77.79$ $64.25$ $51.75$ $53.42$ $30.15$ $1.50$ $20.95$ $9.95$ $8.15$ $3.95$ $25.18$ $35.96$
    Model
    Food101
    CIFAR10
    CIFAR100
    Birdsnap
    SUN397
    Cars
    Aircraft
    VOC2007
    DTD
    Pets
    Caltech101
    Flowers
    MNIST
    StratchCLIP $11.50$ $59.70$ $27.94$ $4.80$ $5.30$ $37.20$ $14.60$ $42.30$ $35.50$ $48.60$ $48.72$ $8.33$ $89.75$
    SpikeCLIP $9.80$ $58.03$ $26.66$ $4.50$ $5.45$ $35.80$ $12.20$ $42.45$ $36.80$ $44.89$ $48.28$ $9.02$ $88.55$
    Appendix F Comparison of Computing Energy Consumption
    According to [65], the theoretical Computing energy consumption of layer $l$ in a SNN can be calculated as:
@@ Appendix F Comparison of Computing Energy Consumption
    where FLOPs is the floating point operations of $b$, which is the number of multiply-and-accumulate (MAC) operations. Assuming that the MAC and AC operations are implemented on the 45nm hardware [68], where $E_{MAC}=4.6pJ$ and $E_{AC}=0.9pJ$ ($1$J $=10^{3}$ mJ $=10^{12}$ PJ).
@@
    Therefore, we estimate the theoretical Computing energy consumption of SpikeCLIP as follows:
@@
    where SNN FC and SNN Conv are the fully connected linear layer and the convolutional layer with neurons in SpikeCLIP respectively. As shown in Equation 10, the SOPs of $m$ SNN Fully Connected Layer (FC), $n$ SNN Convolutional layers are added together and multiplied by $E_{AC}$.
    We refer to [68], assuming that MAC and AC operations are implemented on 45nm hardware (the calculation of power consumption in this hardware only involves MAC and AC operations) since SpikeCLIP and ScratchCLIP have the same architecture except for pulsar neurons, We can calculate the energy consumption reduction (ECR) by equations 7, 8, 9 and 10 as the following expression equation:
@@
    Appendix G Limitations
    In this study, we have embarked on one of the initial endeavors to employ spiking neural networks (SNNs) in multimodal tasks, with a particular emphasis on classification tasks. It would indeed be intriguing to broaden the scope of this work to encompass generative tasks, such as image captioning and visual question answering. We relied on an existing CLIP to pre-train SpikeCLIP by distilling knowledge from the conventional CLIP.
    Another limitation lies in our estimation of energy consumption. While we follow previous studies [65, 39] for calculating energy consumption, their methods only consider computing energy. However, memory access and data movement are also significant factors that affect the overall energy consumption of SNNs, and we further discuss their impact as follows:
    Memory access plays a crucial role in the energy consumption of SNNs, especially when considering the deployment of SNNs on real hardware platforms like FPGAs or specialized neuromorphic chips, such as Loihi [69]. Unlike conventional ANNs, where the energy consumption is primarily driven by the Multiply-and-Accumulate (MAC) operations, SNNs rely heavily on sparse event-driven computations. While the computational load in SNNs may appear lower due to fewer spikes being processed compared to the dense activations in ANNs, the energy cost associated with memory access is often overlooked. In SNNs, the energy expenditure is not just due to MAC or Accumulate (AC) operations but also the need to frequently read and write spike data to memory, especially when dealing with large-scale networks. The memory access patterns in SNNs can result in significant energy overhead, particularly in scenarios where spikes are stored in large buffers or memory arrays. This becomes even more critical when hardware platforms rely on external memory, where the energy cost of fetching and storing spikes can dominate the overall energy budget. Thus, while SNNs might appear more energy-efficient at first glance due to their sparse nature, the energy consumption tied to memory access can significantly reduce these advantages, especially in hardware with high memory access latencies or limited bandwidth. This aspect of energy consumption must be carefully considered when evaluating the overall efficiency of SNNs.